OpenStudy (anonymous):

Solve the problem. Susan purchased some municipal bonds yielding 7% annually and some certificates of deposit yielding 9% annually. If Susan's investment amounts to $19,000 and the annual income is $1590, how much money is invested in bonds and how much is invested in certificates of deposit? a. $13,500 in bonds; $5500 in certificates of deposit b. $5500 in bonds; $13,500 in certificates of deposit c. $13,000 in bonds; $6000 in certificates of deposit d. $6000 in bonds; $13,000 in certificates of deposit

OpenStudy (anonymous):

Hej satellite help me please....

OpenStudy (anonymous):

sorry it should be \[.07x+.09(19000-x)=1590\]

OpenStudy (anonymous):

i get x = 6000

OpenStudy (anonymous):

Solve the problem. Find all values of k so that the given points are \sqrt {29} units apart. (-5, 5), (k, 0) a. 3, 7 b. 7 c. -3, -7 d. -7

hero (hero):

C'mon one question per thread

hero (hero):

I wanted to discuss the first question

OpenStudy (anonymous):

What do you think Hero?

hero (hero):

Satellite, did you set that up as two equations then substitute y into the first one?

hero (hero):

.07x + .09y = 1590 x + y = 19000

hero (hero):

?

OpenStudy (anonymous):

Hero the right answer is not D_= 6000 in bonds, is it?

OpenStudy (anonymous):

no i reasoned as follows. put x = amount in 7% CD, then since the total is 19000 the rest must be 19000 - x

hero (hero):

But the equation comes out in the same manner as what you posted

OpenStudy (anonymous):

yes of course. i put x = amount in 7% and got x = 6000 so that is the amount in municipal bonds. the rest is in the 9% stuff

hero (hero):

What does x and y represent?

OpenStudy (anonymous):

i guess i just did the y = 19000 - x part in my head

hero (hero):

I'm thinking x is the amount invested in bonds and y is the amount invested in certificates, right?

hero (hero):

I mean, the answer must be d

OpenStudy (anonymous):

Ok the same i think...

OpenStudy (anonymous):

Solve the problem. Find all values of k so that the given points are \sqrt {29} units apart. (-5, 5), (k, 0) a. 3, 7 b. 7 c. -3, -7 d. -7